OpenStudy (anonymous):

I need help finding the solution set for this problem. (x+16)(x-18)(x+5)>0 The Solution set is {x|___}

myininaya (myininaya):

first what are the zeros of \[f(x)=(x+16)(x-18)(x+5)\]

myininaya (myininaya):

I will give you a hint there are 3 zeros

OpenStudy (anonymous):

totally lost

myininaya (myininaya):

what x makes f=0 what about x=-16? when you plug that in, isn't the first factor then equal to zero?

myininaya (myininaya):

so that is one zero can you see the other two now?

OpenStudy (anonymous):

I see, so how do I plug the answer {x| ____} in here?

myininaya (myininaya):

we haven't got there

myininaya (myininaya):

so what are the other two zeros?

OpenStudy (anonymous):

18, -5

myininaya (myininaya):

good!

myininaya (myininaya):

-----|-----|-----|----- -16 -5 18 so we need to test these intervals to see if we have f>0 or if f<0

OpenStudy (anonymous):

ok

myininaya (myininaya):

so try pluggin and -18 into f and plug in -10 into f and plug in 0 into f and plug in 20 into f just choose a number in each interval

OpenStudy (anonymous):

honestly I don't know

myininaya (myininaya):

we need to test (-inf,-16) choose a number in this interval and plug into f we need to test (-16,-5) choose a number in this interval and plug into f we need to test (-5,18) choose a number in this interval and plug into f we need to test (18,inf) choose a number in this interval and plug into f

myininaya (myininaya):

for example in that first interval I mention a number in that interval is -18 so I will plug -18 into f recall \[f(x)=(x+16)(x-18)(x+5) \] \[f(-18)=(-18+16)(-18-18)(-18+5)=(-)(-)(-)=(-) <0\]

myininaya (myininaya):

you want to try the next interval

OpenStudy (anonymous):

WOW this is so complicated for me at this time at night. LOL

myininaya (myininaya):

i just replaced the x's with -18 i didn't really care about getting the actually number that that mess equaled i was trying to decide if f(-18) was going to be positive or negative

myininaya (myininaya):

-18+16= negative -18-18=negative -18+5=negative the product of these outcomes is negative right negative numbers are less than 0 so thats why you see <0

OpenStudy (anonymous):

Thank you! I will try tomorrow I honestly can think right now.

myininaya (myininaya):

Ok goodnight!